# Does LLM Focus on the Right Words? Mitigating Context Bias in LLM-based Recommenders

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Recommender systems
@@ 1. Introduction
With remarkable open-domain knowledge and semantic reasoning capabilities (Guo et al., 2025; Achiam et al., 2023), Large Language Models (LLMs) have been extensively explored for integration into recommendation systems (RS) (Wu et al., 2024). One prominent approach involves positioning LLMs as the central recommendation backbone (Li et al., 2023c; Bao et al., 2025, 2023; Zhu et al., 2024; Wang et al., 2025b, 2024b; Liao et al., 2024b; Kim et al., 2024; Hou et al., 2024; Wang and Lim, 2023; Tan et al., 2024; Zheng et al., 2024a). These methods express items as textual descriptions (e.g., titles), and construct language prompts based on users’ past interactions, which are then used to instruct LLMs to predict users’ future interactions. Figure 1 illustrates the mechanism of such LLM-based recommendation. LLMs operate at a fine-grained semantic token level, sequentially generating tokens of the predicted items by analyzing their nuanced semantic relations with the user’s previous interactions and other auxiliary information (e.g., task descriptions, prefix tokens of the prediction). This fine-grained paradigm enables the capture of subtle semantic patterns in user interests and thus represents a promising direction for advancing recommender systems (Bao et al., 2025).
Figure 1. Illustration of LLM‑based recommendations and context bias, wherein the model exhibits an over‑reliance on auxiliary tokens and insufficient utilization of interaction tokens during generation.
@@
To better align LLMs with recommendation objectives and capture collaborative filtering signals, Supervised Fine-Tuning (SFT) is commonly employed (Bao et al., 2025, 2023; Liao et al., 2024b; Zhu et al., 2024). In this strategy, each prompt is paired with the target item description, and the LLM is fine-tuned to generate the correct prediction. This procedure enables the model to capture semantic relationships between prompts and targets present in the training data, often resulting in substantial performance gains.
@@
Figure 3. Distribution of Top‑1 recommended items generated by SFT-trained LLMs across five group defined according to the semantic relevance of items to the auxiliary tokens (Group 1: highest relevance, Group 5: lowest relevance). We also present the distribution of the test set across the same five groups for comparison.
@@ 2.1. LLM-based Recommendation
The remarkable success of LLMs across diverse domains (Wu et al., 2023b; Ouyang et al., 2022; Tang et al., 2024; Chen et al., 2025c; Wang et al., 2025a; Chen et al., 2025b) has spurred growing interest in their application to recommendation systems (RS) (Wu et al., 2024). A prominent approach is to directly leverage LLMs as recommenders (Li et al., 2023c). As shown in Figure 1, this paradigm constructs a language prompt $x=[\,x^{\text{task}};\,x^{\text{user}}\,]$, where $x^{\text{task}}$ represents the task description and $x^{\text{user}}$ denotes the textual form (e.g., titles) of a user’s historical interactions. This prompt then guides the LLM to generate the descriptions of recommended items $y$. Notably, LLMs operate at a fine-grained semantic token level, sequentially generating tokens of the predicted items according to the model estimated probability $P_{\theta}(y_{t}|x,y_{<t})$, where $y_{t}$ denotes the $t$-th predictive tokens and $y_{<t}$ denotes the prefix tokens of the prediction. This fine-grained token-level paradigm has the potential to capture subtle semantic patterns in user preferences.
@@ 2.2. Analyses on Context Bias
In this section, we first identify the context bias in fine-tuning LLM for recommendation, followed by discussing its negative effect. We then analyze the underlying causes of this bias and discuss why existing methods can not effectively address this issue.  [^1] The configuration of the empirical analysis in this section is consistent with our main experimental setup in Section 4.1.3. Additional experiments on other Prompts and LLMs are provided in the Appendix A.1.
@@ 2.2.2. Negative Effect of Context Bias
Here we mask the user historical information in the original prompt template with the placeholder ’N/A’, and compute the probability of the tokens for each item. This measure serves as an indicator of the relevance between the input and output, a practice commonly adopted for estimating semantic relevance (Kauf et al., 2024). Based on this metric, items are sorted and evenly divided into five groups, with Group 1 containing the items of highest relevance and Group 5 the least.
@@ 4. Experiments
  • RQ1: How does GDRT perform compared to existing SOTA methods?
@@
  • RQ4: How do the hyperparameters affect the performance of GDRT?
  • RQ5: How does the efficiency of GDRT compare with baseline methods?
@@ 4.1.1. Datasets
Three widely used real-world datasets—Amazon Toys and Games, Amazon Clothing, Shoes and Jewelry, and Amazon Books [^3] https://jmcauley.ucsd.edu/data/amazon/index_2014.html—are employed in our experiments. These datasets are commonly adopted in studies on LLM-based recommenders (Cui et al., 2024; Bao et al., 2025; Cao et al., 2024; Lee et al., 2024; Li et al., 2023a; Wang et al., 2025b; Lee et al., 2024). To ensure fair comparison, we follow the data preprocessing procedures employed in recent literature (Bao et al., 2025; Cui et al., 2024; Wang et al., 2025b; Bao et al., 2024). Specifically, we first apply the 5-core setting to the raw datasets. For user interaction sequences exceeding 11 interactions, we segment the sequences using a sliding window of length 11. The segmented sequences are then sorted in ascending order by timestamp and partitioned into training, validation, and test sets in an 8:1:1 ratio. Due to the large size of Amazon Books, we randomly retain 100,000 items prior to 5-core processing. The statistics of the processed datasets are summarized in Table 1.
@@ 4.1.2. Baselines
The methods compared fall into several categories:
  • Traditional RS (SASRec (Kang and McAuley, 2018), SASRec++ (Lepage et al., 2025), DROS (Yang et al., 2023)): SASRec employs a self-attention-based model to effectively capture user interests. SASRec++ further improves performance using the softmax loss. DROS integrates DRO to enhance the model’s robustness against distributional shifts.
  • LLM-based RS (SFT (Bao et al., 2025), CFT (Zhang et al., 2024b), MSL (Wang et al., 2025b), LLaRA (Liao et al., 2024b), A-LLM (Kim et al., 2024)): SFT introduces instruction-tuning templates to fine-tune LLMs on RS datasets. CFT incorporates a causal loss to strengthen the behavior sequence modeling capabilities of LLMs. MSL improves the loss function specifically tailoring it to optimize recommendation-oriented objectives. LLaRA incorporates embeddings from traditional models to enhance the ability of LLMs to capture collaborative signals. A-LLM further aligns these embeddings with their corresponding textual information.
  • Debiasing Methods for LLM-based RS (Reweight (Jiang et al., 2024), SPRec (Gao et al., 2025a), D3 (Bao et al., 2024)): Reweight addresses popularity bias by balancing recommendations using pre-calculated item weights. SPRec develops negative sampling techniques based on Direct Preference Optimization (DPO) (Rafailov et al., 2024) to reduce popularity bias. D3 mitigates amplification bias through enhanced decoding strategies.
@@ 4.1.3. Implementation Details
For all LLM-based methods, we adopt LLaMA3.2-3B (Dubey et al., 2024) as the backbone, with the number of training epochs set to 5. The prompt design follows (Bao et al., 2025; Wang et al., 2025b). For inference, we employ Constrained Beam Search (CBS), following prior work (Wang et al., 2025b; Bao et al., 2024), to ensure that the recommended items are drawn from the item set. The beam size is fixed at 10. For evaluation, we evaluate the NDCG@5 on the validation set for each epoch’s checkpoint and select the checkpoint with the highest score for testing. The corresponding results on the test set are reported as the final performance. For the hyperparameters in GDRT, the number of groups $G$ is tuned from $\{2,5,10\}$, and $\tau$ is tuned from $\{0.1,0.2,0.3,0.5,1.0\}$. To ensure fair comparisons, we utilize the source code provided by the original authors and tune the hyperparameters of all baseline methods according to the guidelines specified in their respective publications.
@@ 4.2. Performance Comparison (RQ1 & RQ2)
Table 2 presents a comparative analysis of the proposed GDRT method against the baselines. Overall, GDRT demonstrates substantial performance improvements across all datasets. By mitigating context bias, reducing excessive reliance on auxiliary tokens, and enhancing the model’s ability to capture user-specific behavioral patterns, GDRT delivers marked gains in both accuracy and fairness. Specifically, GDRT achieves an average improvement of 24.29% in NDCG@10 compared with SOTA methods. In terms of fairness, GDRT significantly alleviates the model’s tendency to recommend items strongly correlated with auxiliary information, leading to an average reduction of 37.43% in MGU@5 relative to the SOTA baselines. In contrast, other comparison methods yield only marginal improvements, and their lack of explicit mechanisms to address context bias results in limited fairness enhancement.
Furthermore, we evaluate the performance of GDRT when integrated into other advanced LLM-based RS. As shown in Table 3, GDRT consistently achieves significant performance gains and improves fairness across all methods, indicating that it can be seamlessly incorporated into diverse LLM-based recommenders while exhibiting strong generalization capability.
Figure 8. Distribution of Top‑1 recommended items from GDRT-trained LLMs across five groups defined by auxiliary-token relevance (Group 1: highest relevance, Group 5: lowest relevance), with test-set distribution for comparison.
@@
4.3. In-depth Analysis (RQ3)
In this section, we present an empirical analysis of the effectiveness of GDRT in mitigating context bias. As shown in Figure 7, GDRT considerably reduces the ratio of attribution values assigned to auxiliary tokens relative to interaction tokens, compared with SFT, indicating that GDRT effectively alleviates the problem of over-reliance on auxiliary tokens. Furthermore, GDRT markedly mitigates the unfairness induced by context bias. As shown in Figure 8, the recommendation distributions generated by GDRT more closely align with the item distribution in the test set. In addition, Figure 9 reveals that GDRT alleviates the issue observed in SFT, where the training process continuously amplifies the recommendation of items highly correlated with auxiliary tokens.
@@ 4.5. Efficiency Comparison (RQ5)
This section compares the efficiency and performance of GDRT with baseline methods. As illustrated in Figure 12, GDRT demonstrates both optimal performance and high computational efficiency. In contrast, the DPO-based method SPRec requires an additional reference model and incorporates extra negative samples during training, leading to a considerable increase in runtime. CFT incurs higher computational costs by introducing additional counterfactual samples during training. Although Reweight and D3 incur negligible training overhead, their performance improvements remain limited.
@@ 5.1. Sequential Recommendation
Sequential recommendation focuses on predicting the next item a user will be interested in based on their historical interactions. Compared to collaborative filtering (Chen et al., 2025a, 2024a; Yang et al., 2024; Zhang et al., 2025a; Yang et al., 2025; Wang et al., 2024a), sequential recommendation incorporates temporal information and places greater emphasis on capturing the evolving patterns of user interests. With the advancement of deep learning, numerous architectures based on deep neural networks have been introduced into the sequential recommendation. For example, GRU4Rec (Hidasi et al., 2015) employs RNNs, while Caser (Tang and Wang, 2018) utilizes CNNs to effectively capture long-term dependencies and modeling user interest patterns from historical behavior. More advanced models such as SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019) leverage self-attention mechanisms (Vaswani, 2017), enabling the identification of the most relevant parts within the sequence. Due to the dynamic nature of data distributions as time evolves (Wang et al., 2024c, d), DROS (Yang et al., 2023) introduces DRO (Rahimian and Mehrotra, 2019; Wu et al., 2023a) to further enhance the model’s robustness against distributional shifts caused by temporal changes. The readers may refer to the survey (Fang et al., 2020; Wang et al., 2019) for more details.
@@ 5.2. Biases in LLM-based Recommendation
Large Language Models (LLMs), with their powerful capabilities in comprehension, reasoning, and extensive knowledge (Dubey et al., 2024; Achiam et al., 2023; Guo et al., 2025; Team et al., 2024), have been widely applied to recommendation systems (Wu et al., 2024; Wang et al., 2025c; Cui et al., 2025b, a). One prominent paradigm is LLM-based RS (Li et al., 2023c), which directly leverages LLMs as the backbone of the recommender. Subsequent studies have explored fine-tuning LLMs on domain-specific recommendation datasets to enhance their recommendation capabilities (Li et al., 2023c; Bao et al., 2025, 2023; Zhu et al., 2024; Wang et al., 2025b, 2024b; Liao et al., 2024b; Kim et al., 2024; Chen et al., 2024b).
Recent studies have extensively explored bias and fairness issues in LLM-based RS, such as popularity bias (Gao et al., 2025b; Lu et al., 2025; Jiang et al., 2024; Gao et al., 2025a; Liao et al., 2024a; Lin et al., 2025), position bias (Chen et al., 2023; Ma et al., 2023; Hou et al., 2024; Bito et al., 2025; Jiang et al., 2025; Dai et al., 2024; Luo et al., 2025; Zhang et al., 2024a), amplification bias (Bao et al., 2024), and bias stemming from LLMs’ preferences for specific item attributes (Li et al., 2023b; Deldjoo, 2024; Jiang et al., 2024; Zhang et al., 2025b). Nevertheless, existing research has largely overlooked context bias, which arises during fine-tuning and reflects the inherent over-reliance of LLMs on auxiliary tokens. Since existing debiasing methods fail to account for this factor, their effectiveness remains limited. CFT (Zhang et al., 2024b) seeks to enhance the modeling of users’ historical interactions, but suffers from objective misalignment, weight selection challenges, and high computational cost, restricting its applicability. These limitations are further discussed in Section 2.2.4.

### v3

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792607isbn: 979-8-4007-2307-0/2026/04ccs: Information systems Recommender systems
@@ 1. Introduction
Figure 1. Illustration of LLM‑based recommendations and context bias, wherein the model exhibits an over‑reliance on auxiliary tokens (i.e., Task description, Prefix tokens) and insufficient utilization of User interaction during generation.
@@
With remarkable open-domain knowledge and semantic reasoning capabilities (Guo et al., 2025; Achiam et al., 2023), Large Language Models (LLMs) have been extensively explored for integration into recommendation systems (RS) (Wu et al., 2024). One prominent approach involves positioning LLMs as the central recommendation backbone (Wang et al., 2025b; Li et al., 2023b; Bao et al., 2025, 2023; Zhu et al., 2024; Wang and Lim, 2023; Tan et al., 2024; Zheng et al., 2024a; Wang et al., 2024b; Liao et al., 2024b; Kim et al., 2024). These methods express items as textual descriptions (e.g., titles), and construct language prompts based on users’ past interactions, which are then used to instruct LLMs to predict users’ future interactions. Figure 1 illustrates the mechanism of such LLM-based recommendation. LLMs operate at a fine-grained semantic token level, sequentially generating tokens of the predicted items by analyzing their nuanced semantic relations with the user’s previous interactions and other auxiliary information (e.g., task descriptions, prefix tokens of the predicted item). This fine-grained paradigm enables the capture of subtle semantic patterns in user interests and thus represents a promising direction for advancing recommender systems (Bao et al., 2025).
To better align LLMs with recommendation objectives and capture collaborative filtering signals, Supervised Fine-Tuning (SFT) is commonly employed (Bao et al., 2025, 2023; Liao et al., 2024b; Wang et al., 2025b). In this strategy, each prompt is paired with the target item description, and the LLM is fine-tuned to generate the correct prediction. This procedure enables the model to capture semantic relationships between prompts and targets present in the training data, often resulting in substantial performance gains.
@@
Figure 3. Distribution of Top‑1 recommended items generated by the SFT-trained model across five group defined according to the semantic relevance of items to the auxiliary tokens (Group 1: highest relevance, Group 5: lowest relevance). We also present the distribution of the test set across the same five groups for comparison.
@@ 2.1. LLM-based Recommendation
The remarkable success of LLMs across diverse domains (Wu et al., 2023b; Ouyang et al., 2022; Tang et al., 2024; Chen et al., 2025c; Wang et al., 2025a; Chen et al., 2025b) has spurred growing interest in their application to recommendation systems (RS) (Wu et al., 2024). A prominent approach is to directly leverage LLMs as recommenders (Li et al., 2023b). As shown in Figure 1, this paradigm constructs a language prompt $x=[\,x^{\text{task}};\,x^{\text{user}}\,]$, where $x^{\text{task}}$ represents the task description and $x^{\text{user}}$ denotes the textual form (e.g., titles) of a user’s historical interactions. This prompt then guides the LLM to generate the descriptions of recommended items $y$. Notably, LLMs operate at a fine-grained semantic token level, sequentially generating tokens of the predicted items according to the model estimated probability $P_{\theta}(y_{t}|x,y_{<t})$, where $y_{t}$ denotes the $t$-th predictive tokens and $y_{<t}$ denotes the prefix tokens of the prediction. This fine-grained token-level paradigm has the potential to capture subtle semantic patterns in user preferences.
@@ 2.2. Analyses on Context Bias
In this section, we first identify the context bias in fine-tuning LLM for recommendation, followed by discussing its negative effect. We then analyze the underlying causes of this bias and discuss why existing methods can not effectively address this issue.  [^1] The experimental configuration in this section is consistent with our main experimental setup described in Section 4.1.3. We also provide additional analyses on other Prompt templates and LLMs in Appendix A.1.
@@ 2.2.2. Negative Effect of Context Bias
Here we mask the user historical information in the original prompt template with the placeholder ‘N/A’, and compute the probability of the tokens for each item. This measure serves as an indicator of the relevance between the input and output, a practice commonly adopted for estimating semantic relevance (Kauf et al., 2024). Based on this metric, items are sorted and evenly divided into five groups, with Group 1 containing the items of highest relevance and Group 5 the least.
@@ 4. Experiments
  • RQ1: How does GDRT perform compared to SOTA methods?
@@
  • RQ4: How do the hyperparameters affect the GDRT?
  • RQ5: How does the efficiency of GDRT compare with baselines?
@@ 4.1.1. Datasets
Three widely used real-world datasets—Amazon Toys and Games, Amazon Clothing, Shoes and Jewelry, and Amazon Books [^3] https://jmcauley.ucsd.edu/data/amazon/index_2014.html—are employed in our experiments. To ensure fair comparison, we follow the data preprocessing procedures employed in recent literature (Bao et al., 2025; Cui et al., 2024; Wang et al., 2025b). Specifically, we first apply the 5-core setting to the raw datasets. For user interaction sequences exceeding 11 interactions, we segment the sequences using a sliding window of length 11. The segmented sequences are then sorted in ascending order by timestamp and partitioned into training, validation, and test sets in an 8:1:1 ratio. Due to the large size of Amazon Books, we randomly retain 100,000 items prior to 5-core processing. The statistics of the processed datasets are summarized in Table 1.
@@ 4.1.2. Baselines
The methods compared fall into several categories: (1) Traditional RS: SASRec (Kang and McAuley, 2018), SASRec++ (Lepage et al., 2025), DROS (Yang et al., 2023). (2) LLM-based RS: SFT (Bao et al., 2025), CFT (Zhang et al., 2024b), MSL (Wang et al., 2025b), LLaRA (Liao et al., 2024b), A-LLM (Kim et al., 2024). (3) Debiasing for LLM-based RS: Reweight (Jiang et al., 2024), SPRec (Gao et al., 2025a), D3 (Bao et al., 2024). For a detailed description, see Appendix A.2.
@@ 4.1.3. Implementation Details
For all LLM-based methods, we adopt LLaMA3.2-3B (Dubey et al., 2024) as the backbone, with the number of training epochs set to 5. The prompt design follows (Bao et al., 2025). For inference, we consistently employ Constrained Beam Search (CBS) across all baselines, following prior work (Wang et al., 2025b; Bao et al., 2024), to ensure that the recommended items are drawn from the item set. The beam size is fixed at 10. For evaluation, we evaluate the NDCG@5 on the validation set for each epoch’s checkpoint and select the checkpoint with the highest score for testing. The corresponding results on the test set are reported as the final performance. For the hyperparameters in GDRT, the number of groups $G$ is tuned from $\{2,5,10\}$, and $\tau$ is tuned from $\{0.1,0.2,0.3,0.5,1.0\}$. To ensure fair comparisons, we utilize the source code provided by the original authors and tune the hyperparameters of all baseline methods according to the guidelines specified in their respective publications.
@@ 4.2. Performance Comparison (RQ1 & RQ2)
Table 2 presents a comparative analysis of the proposed GDRT method against the baselines. Overall, GDRT demonstrates substantial performance improvements across all datasets. By mitigating context bias, reducing excessive reliance on auxiliary tokens, and enhancing the model’s ability to capture user-specific behavioral patterns, GDRT delivers marked gains in both accuracy and fairness. Specifically, GDRT achieves an average improvement of 24.29% in NDCG@10 and 37.43% reduction in MGU@5 compared with the best baseline. In contrast, other comparison methods yield only marginal improvements over SFT, and their lack of explicit mechanisms to address context bias results in limited fairness enhancement. Furthermore, since GDRT operates solely by modifying the loss function, it can be easily integrated into existing LLM-based RS without altering their architectures. To verify its general applicability, we further evaluate GDRT when incorporated into several advanced LLM-based recommendation methods. As shown in Table 3, GDRT consistently yields significant improvements in both performance and fairness across all evaluated methods, demonstrating its strong generalization capability. Additional comparisons using alternative prompt templates and LLMs beyond those used in the main experiments are provided in Appendix A.3.
Figure 8. Distribution of Top‑1 recommended items from the GDRT-trained model across five groups defined by auxiliary-token relevance (Group 1: highest relevance, Group 5: lowest relevance), with test-set distribution for comparison.
4.3. In-depth Analysis (RQ3)
In this section, we present an empirical analysis of the effectiveness of GDRT in mitigating context bias. First, we perform FAA on both SFT- and GDRT-trained models, measuring the ratio of attribution values assigned to auxiliary tokens versus interaction tokens. As shown in Figure 7, GDRT yields a significantly lower ratio than SFT, indicating that GDRT effectively alleviates the model’s over-reliance on auxiliary tokens. Furthermore, we analyze the recommendation distributions on different item groups with varying degrees of auxiliary-token relevance. As illustrated in Figure 8, the recommendation distribution generated by the GDRT-trained model closely align with the distribution of the test set. This improvement can be attributed to GDRT’s ability to suppress the tendency of SFT to progressively amplify items highly correlated with auxiliary tokens as shown in Figure 9. Overall, these results demonstrate that GDRT effectively mitigates context bias.
@@ 4.5. Efficiency Comparison (RQ5)
This section compares the efficiency and performance of GDRT with baselines. As illustrated in Figure 12, GDRT demonstrates both optimal performance and high computational efficiency, incurring negligible overhead compared to SFT. In contrast, CFT introduces additional counterfactual samples leading to substantially higher computational costs. SPRec, based on DPO (Rafailov et al., 2024), requires an extra reference model and negative samples during training, resulting in a considerable increase in runtime. Although Reweight and D3 are relatively efficient, their performance improvements are limited.
@@ 5.1. Sequential Recommendation
Sequential recommendation focuses on predicting the next item a user will be interested in based on their historical interactions. Compared to collaborative filtering (Chen et al., 2025a, 2024a; Yang et al., 2024; Zhang et al., 2025a; Yang et al., 2025; Wang et al., 2024a), sequential recommendation incorporates temporal information and places greater emphasis on capturing the evolving patterns of user interests. With the advancement of deep learning, numerous architectures based on deep neural networks have been introduced into the sequential recommendation. For example, GRU4Rec (Hidasi et al., 2015) employs RNNs, while Caser (Tang and Wang, 2018) utilizes CNNs to effectively capture long-term dependencies and modeling user interest patterns from historical behavior. More advanced models such as SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019) leverage self-attention mechanisms (Vaswani, 2017), enabling the identification of the most relevant parts within the sequence. Due to the dynamic nature of data distributions as time evolves (Wang et al., 2024c, d), DROS (Yang et al., 2023) introduces DRO (Rahimian and Mehrotra, 2019; Wu et al., 2023a) to further enhance the model’s robustness against distributional shifts caused by temporal changes. The readers may refer to the survey (Fang et al., 2020) for more details.
@@ 5.2. Biases in LLM-based Recommendation
Large Language Models (LLMs), with their powerful capabilities in comprehension, reasoning, and extensive knowledge (Dubey et al., 2024; Achiam et al., 2023; Guo et al., 2025; Team et al., 2024), have been widely applied to recommendation systems (Wu et al., 2024; Wang et al., 2025c; Cui et al., 2025b, a). One prominent paradigm is LLM-based RS (Li et al., 2023b), which directly leverages LLMs as the backbone of the recommender. Subsequent studies have explored fine-tuning LLMs on domain-specific recommendation datasets to further enhance their recommendation capabilities (Li et al., 2023b; Bao et al., 2025, 2023; Zhu et al., 2024; Wang et al., 2025b, 2024b; Liao et al., 2024b; Kim et al., 2024; Chen et al., 2024b).
Recent studies have extensively explored bias and fairness issues in LLM-based RS, such as popularity bias (Gao et al., 2025b; Lu et al., 2025; Jiang et al., 2024; Gao et al., 2025a; Liao et al., 2024a; Lin et al., 2025), position bias (Chen et al., 2023; Ma et al., 2023; Hou et al., 2024; Bito et al., 2025; Jiang et al., 2025; Dai et al., 2024; Luo et al., 2025; Zhang et al., 2024a), amplification bias (Bao et al., 2024), and bias stemming from LLMs’ preferences for specific item attributes (Li et al., 2023a; Deldjoo, 2024; Jiang et al., 2024; Zhang et al., 2025b). Nevertheless, existing research has largely overlooked context bias, which arises during fine-tuning and reflects the inherent over-reliance of LLMs on auxiliary tokens. Since existing debiasing methods fail to account for this factor, their effectiveness remains limited. CFT (Zhang et al., 2024b) seeks to enhance the modeling of users’ historical interactions, but suffers from objective misalignment, weight selection challenges, and high computational cost, restricting its applicability. These limitations are further discussed in Section 2.2.4.